OpenStudy (anonymous):

if 5/6 o a number is 40, then 66 2/3% of the same number is? guys i need ur detail explanation pleaseeeeeee

OpenStudy (anonymous):

5/6x = 40 x= 240/5 = 48 48 * .6666 = 31.9968

OpenStudy (anonymous):

what did you do to have x = 240/5